# Long-time Freeness in the Kicked Top

### v1

@@ Abstract
I Introduction
The study of quantum chaos and the mechanisms by which isolated quantum systems approach thermal equilibrium has fundamentally reshaped our understanding of quantum theory. Recently, attention has shifted toward more refined probes of chaos, with a particular focus on the role of higher-order correlations in quantum dynamics. This shift is motivated by developments at the intersection of various fields. In quantum information theory, higher-order moments play a crucial role in $n$-designs, ensembles of unitaries that approximate the statistical properties of uniformly random unitaries up to the $n$-th moment Delsarte et al. (1991); Dankert (2005); Dankert et al. (2009). This framework has inspired the study of thermalization through higher moments of state ensembles, such as those generated via time evolution or projected ensembles Cotler et al. (2023); Choi et al. (2023); Ho and Choi (2022); Claeys and Lamacraft (2022); Ippoliti and Ho (2022); Lucas et al. (2023); Bhore et al. (2023); McGinley and Fava (2023); Pilatowsky-Cameo et al. (2023); Mark et al. (2024). New insights into the correlation properties of time-evolved states have also been achieved recently by studies on the statistics of energy eigenstates Dymarsky et al. (2018); Huang (2019); Lu and Grover (2019); Murthy and Srednicki (2019a); Shi et al. (2023); Hahn et al. (2024); Jindal and Hosur (2024).
Concurrently, multi-time correlation functions of physical observables $\hat{A}$ have gained prominence for their importance in higher-order hydrodynamics beyond the linear response regime Doyon and Myers (2020); Myers et al. (2020); Fava et al. (2021); Delacrétaz and Mishra (2024) or for the study of quantum information scrambling, quantified by out-of-time-order correlators (OTOCs) $\braket{\hat{A}(t)\hat{A}\hat{A}(t)\hat{A}}$ Maldacena et al. (2016); Hosur et al. (2016); Xu and Swingle (2022); García-Mata et al. (2023). These correlators, which involve unconventional time orderings between times, are known to provide insights into quantum chaos that go beyond standard two-time correlators. For example, in the case of underlying chaotic semi-classical dynamics, the OTOCs encode the quantum Lyapunov exponent at short intermediate times Larkin and Ovchinnikov (1969); Kitaev (20145). To address properties among $2n$ times, the so-called $2n$-OTOCs
@@ I Introduction
have been intensively studied Roberts and Yoshida (2017); Cotler et al. (2017a); Tsuji et al. (2018); Cotler and Hunter-Jones (2020); Leone et al. (2021). These account for $n$-designs information Roberts and Yoshida (2017), and may encode the generalized quantum Lyapunov exponents Pappalardi and Kurchan (2023); Trunin (2023a, b).
@@
The established framework for understanding quantum dynamics of observables $\hat{A}$ is the Eigenstate Thermalization Hypothesis (ETH) Srednicki (1994, 1999); D’Alessio et al. (2016). ETH assumes that matrix elements of $\hat{A}$ in the energy eigenbasis look like pseudorandom matrices with smooth statistical properties. To describe multi-time correlation functions such as in Eq. (1), a complete version of ETH has been introduced, named full ETH, which encompasses correlations among the matrix elements Foini and Kurchan (2019a). The significance of these correlations has attracted substantial interest across fields Prosen (1999); Sonner and Vielma (2017); Foini and Kurchan (2019b); Chan et al. (2019); Murthy and Srednicki (2019b); Richter et al. (2020); Wang et al. (2021); Brenes et al. (2021); Dymarsky (2022); Nussinov and Chakrabarty (2022); Jafferis et al. (2022a, b); Wang et al. (2024); Fava et al. (2023); Pappalardi et al. (2023, 2024). A recent development in this area has been the formalization of full ETH using Free Probability Theory Pappalardi et al. (2022). Free probability is an extension of traditional probability theory to non-commuting variables Voiculescu (1991); Voiculescu et al. (1992); Speicher (1997), which introduces powerful tools for understanding the statistical properties of large matrices, such as those that arise in chaotic quantum systems. Originating in planar field theory Brézin et al. (1978); Cvitanović (1981); Cvitanović et al. (1982), free probability has also played a role in describing correlations in other branches of many-body physics, such as quantum information theory Collins and Nechita (2016), tensor networks Collins and Nechita (2010); Kudler-Flam et al. (2021, 2022); Cheng et al. (2024), disordered systems Movassagh and Edelman (2010, 2011); Chen et al. (2012); Hruza and Bernard (2023); Bauer et al. (2023); Bernard and Hruza (2023), and gravity Berkooz et al. (2019); Penington et al. (2022); Wang (2023); Wu (2024); Chandrasekaran et al. (2023).
@@
The connection between ETH and free probability provides a mathematical structure to describe correlations through non-crossing partitions, and the so-called free cumulants, connected correlation functions that generalize classical cumulants. In particular, free cumulants between observables at $2n$ times, i.e. $\kappa_{2n}(t,0,\dots,t,0)$, provide a means to systematically study the $2n$-OTOC in Eq. (1). This approach leads to the proposal that chaotic dynamics shall result at long times in “free independence” or freeness of observables, which suggests that observables become statistically independent in a specific, non-classical sense as the system evolves over time Cipolloni et al. (2022); Fava et al. (2023). While these ideas are mathematically appealing, their application has been primarily explored in systems evolving with random Wigner matrices Cipolloni et al. (2022) or in fine-tuned models, such as dual-unitary circuits Chen and Kudler-Flam (2024). The physical implications of long-time freeness in realistic chaotic systems remain largely unexplored.
@@
In this work, we address the emergence of freeness at long times by investigating the kicked top, a paradigmatic semiclassical model of quantum chaos Haake et al. (1987); Kuś et al. (1987). The kicked top describes a system of spins driven periodically in time via collective interactions. Thanks to the latter, the system has a well-defined classical limit, corresponding to the thermodynamic limit, and the dimension of the Hilbert space scales only linearly with the system size, which makes it amenable to exact numerics. Furthermore, this model exhibits a transition from integrability to chaos, making it an ideal setting for investigating the interplay between chaotic dynamics, freeness and its breakdown. Despite its simplicity, the kicked top captures many essential features of chaotic quantum systems, allowing us to identify several features that apply to many-body non-integrable systems.
@@
  • Large Deviation Theory of Freeness For systems with exponentially decaying free cumulants, we develop a large deviation approach to freeness, inspired by standard approaches to multifractality in turbulence Benzi et al. (1984); Paladin and Vulpiani (1987), chaos theory Paladin and Vulpiani (1986); Crisanti et al. (1988), and wavefunction localizations Evers and Mirlin (2000); Macé et al. (2019); Sierant and Turkeshi (2022). When the $2n$-OTOCs decay exponentially, one can define a “freeness time-scale” through the quantities $\tau_{n}$, obtained as
@@
The rest of the paper is structured as follows. In Section II we recall the model of the kicked top and we introduce the main concepts of Free Probability - free cumulants and freeness - connected to the full Eigenstate Thermalization Hypothesis. In Section III, we focus on the dynamics of free cumulants; we observe the emergence of long-time freeness only in the chaotic regime and we highlight the relevance of free cumulants for describing specific correlators like the $2n$-OTOCs and the square-commutator. In Section IV, we present the definition of a time scale for the onset of freeness, developing a large deviation theory for freeness in Section IV.1. Finally, in Section IV.2, we study numerically the freeness time-scale, showing freeness multifractality in the model analyzed. At last, Section V is devoted to the conclusions and the open questions.
II Modeling Chaos: Kicked Top,
@@
In this section, we recall the definition of the quantum kicked top and we provide a pedagogical review of the Free Probability approach to the full Eigenstate Thermalization Hypothesis, which we will examine in the subsequent analysis. Readers who are familiar with these topics may proceed directly to our results in Sec. III and Sec. IV.
II.1 Kicked Top
Figure 1: Transition between regular and chaotic regimes for the classical kicked top. The classical variable is the rescaled angular momentum $(X,Y,Z)\equiv 2{(J_{x},J_{y},J_{z})}/{N}$ which lies on a unitary sphere. Here 253 random initial conditions with $Y_{0}>0$ are evolved for 250 kicks through the classical equations of motionII.1.
@@ II.1 Kicked Top
By increasing the kicking strength $\gamma$, the system undergoes a transition between regular and chaotic dynamics: the periodic drive introduces energy into the system, making it chaotic. See Fig. 1 for an illustration of the classical phase-space for different $\gamma$. Being a paradigmatic model of quantum chaos, the classical Haake et al. (1987); Kuś et al. (1987) and quantum Chaudhury et al. (2009); Wang and Robnik (2021) features of this transition have been extensively studied, including its OTOC dynamics Pappalardi et al. (2018); Seshadri et al. (2018); Sieberer et al. (2019); Pilatowsky-Cameo et al. (2020); Lerose and Pappalardi (2020).
@@
II.2 Free Probability in chaotic dynamics
@@
where $D$ is the dimension of the Hilbert space, and focuses on the quasi-energy eigenbasis. For discussions involving the energy eigenstates, as well as canonical and microcanonical averages, we refer the reader to Refs. Pappalardi et al. (2022, 2023); Fava et al. (2023); Pappalardi et al. (2024).
II.2.1 Full Eigenstate Thermalization Hypothesis
According to the Eigenstate Thermalization Hypothesis Srednicki (1994, 1999); D’Alessio et al. (2016), the matrix elements $A_{\alpha\beta}=\bra{\nu_{\alpha}}\hat{A}\ket{\nu_{\beta}}$ of a physical observable $\hat{A}$ in the quasi-energy eigenbasis behave as pseudorandom matrix with smooth statistical properties. The full ETH Foini and Kurchan (2019a) extends the standard formulation to multiple-point correlations between matrix elements. Specifically, statistical averages111 Although matrix elements are completely determined by the eigenvectors of a fixed Hamiltonian, they behave pseudo-randomly, and the statistical average shall be intended over a “fictitious ensemble” that can be defined heuristically in several ways, such as averaging over a small window of energies or adding a small disordered term to the Hamiltonian and averaging over it, see e.g. Deutsch (1991); Foini and Kurchan (2019a). In practice, the average shall be valid “typically” and ETH can be understood as a self-averaging property of matrix elements. of products with distinct indices $\alpha_{1}\neq\alpha_{2}\neq\dots\neq\alpha_{n}$ read
$\overline{A_{\alpha_{1}\alpha_{2}}A_{\alpha_{2}\alpha_{3}}\dots A_{\alpha_{n}\alpha_{1}}}$
$=D^{1-n}\,$ $F^{(n)}\left(\omega_{\alpha_{1}\alpha_{2}},\dots,\omega_{\alpha_{n-1}\alpha_{n}}\right)$ (4a)
where $F^{(n)}$ are some smooth functions of the eigenenergies differences $\omega_{\alpha_{1}\alpha_{2}}=\nu_{\alpha_{1}}-\nu_{\alpha_{2}}$ that define the observable. For $n=1,2$ one recovers the standard ETH Srednicki (1999), where in the Floquet basis $F^{(1)}=\langle\hat{A}\rangle$ is the constant equilibrium average and $F^{(2)}(\omega)=|f(\omega)|^{2}$ is the dynamical correlation function. When indices are repeated, the average of the products factorizes at the leading order in $D$ as
@@ II.2.1 Full Eigenstate Thermalization Hypothesis
As a result of the smoothness of the matrix elements and of the scaling with $D$ in Eq. (4), Ref. Foini and Kurchan (2019a) showed that the correlation functions at each $n$ are organized in terms of sums over distinct indices as
@@
These are referred to as ETH simple loops because, in the diagrammatic representations of the full ETH, they correspond to diagrams where energy indices appear as dots arranged in a single loop. For completeness, a detailed discussion of ETH diagrams and the derivation of this result is provided in App. A. Furthermore, according to Eq. (4), the simple loops can be written as
@@
II.2.2 Free Cumulants
The simple loops $\mathcal{K}_{n}$ in Eq. (6) appearing in ETH description of multi-time correlators can actually be identified as free cumulants, a familiar object in the theory of Free Probability Speicher (1997), see Refs. Speicher (2016); Xia (2019) for an introduction.
Free cumulants $\kappa_{n}$ are connected correlation functions defined from moments of $n$ variables. Let us here consider the case of a time-dependent observable at different times: $\hat{A}(t_{i})$ $i=1,\dots,n$. Free cumulants are defined implicitly and recursively through the moments-free cumulant formula Speicher (1997):
@@ II.2.2 Free Cumulants
While Eq. (7) is just an implicit definition of free cumulants in terms of moments, when ETH applies, the free cumulants correspond to the ETH simple loops $\mathcal{K}_{n}$ defined in Eq. (6) namely Pappalardi et al. (2022)
@@ II.2.3 Freeness
where “asymptotically” indicates that this condition holds as the matrix dimension $D$ diverges. When $A$ and $B$ are free, their mixed moments can be computed from their individual moments only  [^3] For example, consider the implicit definition in Eq. (7) for alternating free variables $A,B$: $\langle AB\rangle$ $=\kappa_{1}(A)\kappa_{1}(B)+\kappa_{2}(A,B)\qquad\text{with}\qquad\kappa_{2}(A,B)=0$ $\to\langle AB\rangle$ $=\langle A\rangle\langle B\rangle\ ,$ $\langle ABAB\rangle=$ $\kappa_{4}(A,B,A,B)+\kappa^{2}_{1}(A)\kappa^{2}_{1}(B)+4\kappa_{1}(A)\kappa_{1}(B)\kappa_{2}(A,B)$ $\quad+\kappa^{2}_{1}(A)\kappa_{2}(B)+\kappa^{2}_{1}(B)\kappa_{2}(A)$ $\quad+2\kappa_{1}(A)\kappa_{3}(B,A,B)+2\kappa_{1}(B)\kappa_{3}(A,B,A)$ $\quad+2\kappa^{2}_{2}(A,B)\qquad\text{with}\qquad\kappa_{n}(A,B,\dots)=0$ $\to\langle ABAB\rangle=$ $\left(\langle A^{2}\rangle-\braket{A}^{2}\right)\braket{B}^{2}+\braket{A}^{2}\braket{B^{2}}\ .$ The mixed moments admit a compact expression given by Mingo and Speicher (2017)
@@
If the vanishing of free cumulants in Eq. (10) holds only up to a certain order $n$, then $A$ and $B$ are termed asymptotically $n$-free. Important examples of free variables can be found in random matrix theory Voiculescu (1991); Mingo and Speicher (2017), specifically when considering randomly rotated deterministic matrices $B$ and $A^{U}=U^{\dagger}AU$. When $U$ is sampled from the Haar ensemble, $A^{U}$ and $B$ are asymptotically free Mingo and Speicher (2017). If $U$ is sampled from an $n$-design (an ensemble of unitaries that approximates the Haar measure up to order $n$), then $A^{U}$ and $B$ are asymptotically $n$-free Fava et al. (2021).
Notably, chaotic time evolution leads to arbitrary asymptotically $n$-freeness of physical observables at long-times, in the limit of diverging system sizes. While this holds for random matrices Cipolloni et al. (2022), it can also be derived for generic many-body Hamiltonians as a consequence of the Eigenstate Thermalization Hypothesis Fava et al. (2023). For quantum observables $\hat{A}(t)$ and $\hat{B}$ satisfying ETH, the infinite-time average of their mixed free cumulants vanishes:
@@
This result is derived using i) the ETH, which expresses free cumulants as sums of simple loops, and ii) a no-resonance condition between eigenvalues [^5] The ETH free cumulant in Eq. (9) is given by sums of simple loops over distinct indices. Hence its time-average selects only terms satisfying $\sum_{i=1}^{m}E_{\alpha_{i}}=\sum_{i=1}^{m}E_{\beta_{i}}$, where $m$ indicates the number of times that $\hat{A}(t)$ and $\hat{B}$ are alternating. The non-resonance condition implies that such equality is satisfied only if the set $\{\alpha_{i}\}$ is a permutation of the set $\{\beta_{i}\}$. Since the sum is only over distinct indices, this implies that the time average vanishes., believed to hold generically in non-integrable systems Srednicki (1999); Linden et al. (2009). This suggests that chaotic time evolution at infinite times is as effective as an $n$-design in inducing freeness among observables.
@@
Long-time freeness implies that mixed moments at late times can be computed using established properties of Free Probability Mingo and Speicher (2017) such as the factorization of Eq. (11). This can be directly applied to out-of-time order correlators. For instance, the OTOC between $\hat{A}(t)$ and $\hat{B}$ at long-times $t\gg 1$ is given by:
@@
$\quad-5\kappa^{3}_{2}(t,0)+3\kappa^{2}_{3}(t,0,t)$ (15e)
$\quad-6\kappa_{2}(t,0)\kappa_{4}(t,0,t,0)\ ,$
@@
III Free cumulants dynamics

### v2

@@ Abstract
1 Introduction
The study of quantum chaos and the mechanisms by which isolated quantum systems approach thermal equilibrium has fundamentally reshaped our understanding of quantum theory. Recently, attention has shifted toward more refined probes of chaos, with a particular focus on the role of higher-order correlations in quantum dynamics. This shift is motivated by developments at the intersection of various fields. In quantum information theory, higher-order moments play a crucial role in $n$-designs, ensembles of unitaries that approximate the statistical properties of uniformly random unitaries up to the $n$-th moment [1, 2, 3]. This framework has inspired the study of thermalization through higher moments of state ensembles, such as those generated via time evolution or projected ensembles [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. New insights into the correlation properties of time-evolved states have also been achieved recently by studies on the statistics of energy eigenstates [14, 15, 16, 17, 18, 19, 20].
Concurrently, multi-time correlation functions of physical observables $\hat{A}$ have gained prominence for their importance in higher-order hydrodynamics beyond the linear response regime [21, 22, 23, 24, 25] or for the study of quantum information scrambling, quantified by out-of-time-order correlators (OTOCs) $\braket{\hat{A}(t)\hat{A}\hat{A}(t)\hat{A}}$ [26, 27, 28, 29]. These correlators, which involve unconventional time orderings between times, are known to provide insights into quantum chaos that go beyond standard two-time correlators. For example, in the case of underlying chaotic semi-classical dynamics, the OTOCs encode the quantum Lyapunov exponent at short intermediate times [30, 31]. To address properties among $2n$ times, the so-called $2n$-OTOCs
@@ 1 Introduction
have been intensively studied [32, 33, 34, 35, 36]. These account for $n$-designs information [32], and may encode the generalized quantum Lyapunov exponents [37, 38, 39].
@@
The established framework for understanding quantum dynamics of observables $\hat{A}$ is the Eigenstate Thermalization Hypothesis (ETH) [40, 41, 42]. ETH assumes that matrix elements of $\hat{A}$ in the energy eigenbasis look like pseudorandom matrices with smooth statistical properties. To describe multi-time correlation functions such as in Eq. (1), a complete version of ETH has been introduced, named full ETH, which encompasses correlations among the matrix elements [43]. The significance of these correlations has attracted substantial interest across fields [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. A recent development in this area has been the formalization of full ETH using Free Probability Theory [61]. Free probability is an extension of traditional probability theory to non-commuting variables [62, 63, 64], which introduces powerful tools for understanding the statistical properties of large matrices, such as those that arise in chaotic quantum systems. Originating in planar field theory [65, 66, 67], free probability has also played a role in describing correlations in other branches of many-body physics, such as quantum information theory [68], tensor networks [69, 70, 71, 72], disordered systems [73, 74, 75, 76, 77, 78], and gravity [79, 80, 81, 82, 83].
@@
The connection between ETH and free probability provides a mathematical structure to describe correlations through non-crossing partitions, and the so-called free cumulants, connected correlation functions that generalize classical cumulants. In particular, free cumulants between observables at $2n$ times, i.e. $\kappa_{2n}(t,0,\dots,t,0)$, provide a means to systematically study the $2n$-OTOC in Eq. (1). This approach leads to the proposal that chaotic dynamics shall result at long times in “free independence” or freeness of observables, which suggests that observables become statistically independent in a specific, non-classical sense as the system evolves over time [84, 57]. While these ideas are mathematically appealing, their application has been primarily explored in systems evolving with random Wigner matrices [84] or in fine-tuned models, such as dual-unitary circuits [85]. The physical implications of long-time freeness in realistic chaotic systems remain largely unexplored.
@@
In this work, we address the emergence of freeness at long times by investigating the kicked top, a paradigmatic semiclassical model of quantum chaos [86, 87]. The kicked top describes a system of spins driven periodically in time via collective interactions. Thanks to the latter, the system has a well-defined classical limit, corresponding to the thermodynamic limit, and the dimension of the Hilbert space scales only linearly with the system size, which makes it amenable to exact numerics. Furthermore, this model exhibits a transition from integrability to chaos, making it an ideal setting for investigating the interplay between chaotic dynamics, freeness and its breakdown. Despite its simplicity, the kicked top captures many essential features of chaotic quantum systems, allowing us to identify several features that apply to many-body non-integrable systems.
@@
  • Large Deviation Theory of Freeness For systems with exponentially decaying free cumulants, we develop a large deviation approach to freeness, inspired by standard approaches to multifractality in turbulence [88, 89], chaos theory [90, 91], and wavefunction localizations [92, 93, 94]. When the $2n$-OTOCs decay exponentially, one can define a “freeness time-scale” through the quantities $\tau_{n}$, obtained as
@@
The rest of the paper is structured as follows. In Section 2 we recall the model of the kicked top and we introduce the main concepts of Free Probability - free cumulants and freeness - connected to the full Eigenstate Thermalization Hypothesis. In Section 3, we focus on the dynamics of free cumulants; we observe the emergence of long-time freeness only in the chaotic regime and we highlight the relevance of free cumulants for describing specific correlators like the $2n$-OTOCs and the square-commutator. In Section 4, we present the definition of a time scale for the onset of freeness, developing a large deviation theory for freeness in Section 4.1. Finally, in Section 4.2, we study numerically the freeness time-scale, showing freeness multifractality in the model analyzed. At last, Section 5 is devoted to the conclusions and the open questions.
2 Modeling Chaos: Kicked Top,
@@
In this section, we recall the definition of the quantum kicked top and we provide a pedagogical review of the Free Probability approach to the full Eigenstate Thermalization Hypothesis, which we will examine in the subsequent analysis. Readers who are familiar with these topics may proceed directly to our results in Sec. 3 and Sec. 4.
2.1 Kicked Top
Figure 1: Transition between regular and chaotic regimes for the classical kicked top. The classical variable is the rescaled angular momentum $(X,Y,Z)\equiv 2{(J_{x},J_{y},J_{z})}/{N}$ which lies on a unitary sphere. Here 253 random initial conditions with $Y_{0}>0$ are evolved for 250 kicks through the classical equations of motion2.1.
@@ 2.1 Kicked Top
By increasing the kicking strength $\gamma$, the system undergoes a transition between regular and chaotic dynamics: the periodic drive introduces energy into the system, making it chaotic. See Fig. 1 for an illustration of the classical phase-space for different $\gamma$. Being a paradigmatic model of quantum chaos, the classical [86, 87] and quantum [95, 96] features of this transition have been extensively studied, including its OTOC dynamics [97, 98, 99, 100, 101].
@@
2.2 Free Probability in chaotic dynamics
@@
where $D$ is the dimension of the Hilbert space, and focuses on the quasi-energy eigenbasis. For discussions involving the energy eigenstates, as well as canonical and microcanonical averages, we refer the reader to Refs. [61, 58, 59, 60].
2.2.1 Full Eigenstate Thermalization Hypothesis
According to the Eigenstate Thermalization Hypothesis [40, 41, 42], the matrix elements $A_{\alpha\beta}=\bra{\nu_{\alpha}}\hat{A}\ket{\nu_{\beta}}$ of a physical observable $\hat{A}$ in the quasi-energy eigenbasis behave as pseudorandom matrix with smooth statistical properties. The full ETH [43] extends the standard formulation to multiple-point correlations between matrix elements. Specifically, statistical averages111 Although matrix elements are completely determined by the eigenvectors of a fixed Hamiltonian, they behave pseudo-randomly, and the statistical average shall be intended over a “fictitious ensemble” that can be defined heuristically in several ways, such as averaging over a small window of energies or adding a small disordered term to the Hamiltonian and averaging over it, see e.g. [102, 43]. In practice, the average shall be valid “typically” and ETH can be understood as a self-averaging property of matrix elements. of products with distinct indices $\alpha_{1}\neq\alpha_{2}\neq\dots\neq\alpha_{n}$ read
$\overline{A_{\alpha_{1}\alpha_{2}}A_{\alpha_{2}\alpha_{3}}\dots A_{\alpha_{n}\alpha_{1}}}$ (4a)
$=D^{1-n}\,$ $F^{(n)}\left(\omega_{\alpha_{1}\alpha_{2}},\dots,\omega_{\alpha_{n-1}\alpha_{n}}\right)$
where $F^{(n)}$ are some smooth functions of the eigenenergies differences $\omega_{\alpha_{1}\alpha_{2}}=\nu_{\alpha_{1}}-\nu_{\alpha_{2}}$ that define the observable. For $n=1,2$ one recovers the standard ETH [41], where in the Floquet basis $F^{(1)}=\langle\hat{A}\rangle$ is the constant equilibrium average and $F^{(2)}(\omega)=|f(\omega)|^{2}$ is the dynamical correlation function. When indices are repeated, the average of the products factorizes at the leading order in $D$ as
@@ 2.2.1 Full Eigenstate Thermalization Hypothesis
As a result of the smoothness of the matrix elements and of the scaling with $D$ in Eq. (4), Ref. [43] showed that the correlation functions at each $n$ are organized in terms of sums over distinct indices as
@@
These are referred to as ETH simple loops because, in the diagrammatic representations of the full ETH, they correspond to diagrams where energy indices appear as dots arranged in a single loop. For completeness, a detailed discussion of ETH diagrams and the derivation of this result is provided in App. A. Furthermore, according to Eq. (4a), the simple loops can be written as
@@
2.2.2 Free Cumulants
The simple loops $\mathcal{K}_{n}$ in Eq. (6) appearing in ETH description of multi-time correlators can actually be identified as free cumulants, a familiar object in the theory of Free Probability [64], see Refs. [103, 104] for an introduction.
Free cumulants $\kappa_{n}$ are connected correlation functions defined from moments of $n$ variables. Let us here consider the case of a time-dependent observable at different times: $\hat{A}(t_{i})$ $i=1,\dots,n$. Free cumulants are defined implicitly and recursively through the moments-free cumulant formula [64]:
@@ 2.2.2 Free Cumulants
While Eq. (7) is just an implicit definition of free cumulants in terms of moments, when ETH applies, the free cumulants correspond to the ETH simple loops $\mathcal{K}_{n}$ defined in Eq. (6) namely [61]
@@ 2.2.3 Freeness
where “asymptotically” indicates that this condition holds as the matrix dimension $D$ diverges. When $A$ and $B$ are free, their mixed moments can be computed from their individual moments only  [^3] For example, consider the implicit definition in Eq. (7) for alternating free variables $A,B$: $\langle AB\rangle$ $=\kappa_{1}(A)\kappa_{1}(B)+\kappa_{2}(A,B)$ $\text{with}\qquad\kappa_{2}(A,B)=0$ $\to\langle AB\rangle$ $=\langle A\rangle\langle B\rangle\ ,$ $\langle ABAB\rangle$ $=\kappa_{4}(A,B,A,B)+\kappa^{2}_{1}(A)\kappa^{2}_{1}(B)$ $+4\kappa_{1}(A)\kappa_{1}(B)\kappa_{2}(A,B)$ $+\kappa^{2}_{1}(A)\kappa_{2}(B)+\kappa^{2}_{1}(B)\kappa_{2}(A)$ $+2\kappa_{1}(A)\kappa_{3}(B,A,B)+2\kappa_{1}(B)\kappa_{3}(A,B,A)$ $+2\kappa^{2}_{2}(A,B)\quad\text{with}\quad\kappa_{n}(A,B,\dots)=0$ $\to\langle ABAB\rangle$ $=\left(\langle A^{2}\rangle-\braket{A}^{2}\right)\braket{B}^{2}+\braket{A}^{2}\braket{B^{2}}\ .$ The mixed moments admit a compact expression given by [105]
@@
If the vanishing of free cumulants in Eq. (10) holds only up to a certain order $n$, then $A$ and $B$ are termed asymptotically $n$-free. Important examples of free variables can be found in random matrix theory [62, 105], specifically when considering randomly rotated deterministic matrices $B$ and $A^{U}=U^{\dagger}AU$. When $U$ is sampled from the Haar ensemble, $A^{U}$ and $B$ are asymptotically free [105]. If $U$ is sampled from an $n$-design (an ensemble of unitaries that approximates the Haar measure up to order $n$), then $A^{U}$ and $B$ are asymptotically $n$-free [23].
Notably, chaotic time evolution leads to arbitrary asymptotically $n$-freeness of physical observables at long-times, in the limit of diverging system sizes. While this holds for random matrices [84], it can also be derived for generic many-body Hamiltonians as a consequence of the Eigenstate Thermalization Hypothesis [57]. For quantum observables $\hat{A}(t)$ and $\hat{B}$ satisfying ETH, the infinite-time average of their mixed free cumulants vanishes:
@@
This result is derived using i) the ETH, which expresses free cumulants as sums of simple loops, and ii) a no-resonance condition between eigenvalues [^5] The ETH free cumulant in Eq. (9) is given by sums of simple loops over distinct indices. Hence its time-average selects only terms satisfying $\sum_{i=1}^{m}E_{\alpha_{i}}=\sum_{i=1}^{m}E_{\beta_{i}}$, where $m$ indicates the number of times that $\hat{A}(t)$ and $\hat{B}$ are alternating. The non-resonance condition implies that such equality is satisfied only if the set $\{\alpha_{i}\}$ is a permutation of the set $\{\beta_{i}\}$. Since the sum is only over distinct indices, this implies that the time average vanishes., believed to hold generically in non-integrable systems [41, 106]. This suggests that chaotic time evolution at infinite times is as effective as an $n$-design in inducing freeness among observables.
@@
Long-time freeness implies that mixed moments at late times can be computed using established properties of Free Probability [105] such as the factorization of Eq. (11). This can be directly applied to out-of-time order correlators. For instance, the OTOC between $\hat{A}(t)$ and $\hat{B}$ at long-times $t\gg 1$ is given by:
@@
$\quad-5\kappa^{3}_{2}(t,0)+3\kappa^{2}_{3}(t,0,t)$
$\quad-6\kappa_{2}(t,0)\kappa_{4}(t,0,t,0)\ ,$ (15e)
@@
3 Free cumulants dynamics